OpenStudy (anonymous):

\[(\log_{x} 3x)(\log_{3}x ^{2})=(\sin1/\cos1)(\sin2/\cos2)(\sin3/\cos3)...(\sin89/\cos89)\]

OpenStudy (jamesj):

First remember that sin(x) = cos(90-x). This should enable you to radically simplify the RHS of this equation

OpenStudy (anonymous):

so should it be (cos(90-x)/cos1)? I'm not sure exactly what to do...

OpenStudy (jamesj):

For example, using that identity, sin 1 = cos 89. Hence that pair of terms cancel. Now what about all of the other terms on the RHS (right hand side)?

OpenStudy (anonymous):

I got the first part being (log3x/logx)\[(2\log_{3}x/\log3)\]

OpenStudy (anonymous):

therefore the (sin1/cos1)=(cos89/cos1)=cos89? am i understanding that correctly?

OpenStudy (jamesj):

sin 1 = cos 89. Hence sin 1 / cos 89 = 1. Every other pair of terms cancel out as well: i.e., sin 2 / cos 88 = 1 sin 3 / cos 87 = 1 ...

OpenStudy (anonymous):

so would the next one be (cos88)?

OpenStudy (jamesj):

next what? Do you understand what I just wrote above?

OpenStudy (anonymous):

not completely. aren't you replacing sin1=cos89 therefore wouldn't it look something like \[\cos 89/\cos 1 => \cos 89\]

OpenStudy (jamesj):

(sin1/cos1)(sin2/cos2)(sin3/cos3)...(sin89/cos89) = (sin1/cos89)(sin2/cos88)....(sin88/cos2)(sin89/cos1) by just rearranging terms. Now each of these fractions is equal to exactly 1.

OpenStudy (anonymous):

what about the first part of the problem with the logs? can you help me understand what to do next?

OpenStudy (jamesj):

I would change base to make the logs all be the same. Here I am changing them all to base \( e\) and hence writing \( \ln = \log_e \)\[ (\log_x 3x)(\log_3 x^2) = \frac{ \ln 3x}{\ln x} \frac{\ln x^2}{\ln 3}\] Now simplify this expression.

OpenStudy (anonymous):

i don't know how. :(

OpenStudy (anonymous):

this was the farthest I've solved it: (log3x/logx)(2log3x/log3)

OpenStudy (jamesj):

Here's one more hint then: \( \ln x^2 = 2 \ln x \). Now notice you can do some cancelation.

OpenStudy (anonymous):

so it comes out to 2x